7) The average annual tuition and fees at all 4-year institutions in the US in 2012 was $26,390 and in 2018 is estimated to be S 35,980. Let y be the average tuition and fees in the year x where x-0 represents the year 201o. a) Write a linear equation (point slope form), which would models the growth in average tuition and fees at all 4-year institutions in the US in terms of the year x. Plot the graph usinga calculator and creating a table by hand if no calculator is used. b) Explain what the slope of this line means in the context of the problem using the calculations and the graph obtained cUse this equation to predict the average tuition and fees at 4-year institutions in the US in the year 2028.
Answer #1

1) we have two points ( 2 , 26390 ) and ( 8, 35980 )

a) slope = ( 35980 - 26390 ) / ( 8 - 2)

= 4795/3

point slope form is

y - 26390 = 4795/3 ( x - 2 )

b) slope means average tuition fees per year

c) in the year 2028

y - 26390 = 4795/3 ( 28- 2)

average tuition and fees in 2028 is

y = $ 67946.67